OpenStudy (anonymous):

Work-Out Corner has 5 more than 3 times as many exercise bicycles as The Gym. Together they have 21 bicycles. Solve the equation x + 3x + 5 = 21 to find the number of bicycles at Work-Out Corner. A. 4 bicycles B. 17 bicycles C. 7 bicycles D. 25 bicycles

OpenStudy (anonymous):

First combine the x's to get one number for them

OpenStudy (anonymous):

#1 is A

OpenStudy (anonymous):

I get error when I look at the graph

OpenStudy (anonymous):

4x+5=21 -5 -5 4x =16 /4 =/4 x =1

OpenStudy (anonymous):

i put A but i got it wrong

OpenStudy (anonymous):

It should be correct (4)+3(4)+5 4+12+5 16+5 21

OpenStudy (anonymous):

4 + 12 (3 x 4) + 5 is 21...its A

OpenStudy (anonymous):

i know, but it marked it wrong :/

OpenStudy (anonymous):

its 17. we didn't finish the problem. 4 is the number at the gym but you have to get 5 more than 4 times 3

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

Another example of why its important to read the whole question